# Machine Translation in the Wild: User Reaction to Xiaohongshu’s Built-In Translation Feature

### 16 Mar 2026 v1

@@ Abstract
The growing integration of machine translation (MT) into social media platforms is transforming how users interact with each other across cultural and linguistic boundaries. This paper examines user reactions to the launch of Xiaohongshu’s built-in translation feature in January 2025. Drawing on a dataset of 6,723 comments collected from 11 official posts promoting the translation function, this paper combines sentiment analysis with thematic analysis to investigate how users perceived and experimented with the function. Results show that reactions were generally positive, particularly for translating posts and comments, although concerns regarding functionality, accessibility, and translation accuracy were also expressed. In addition to evaluative feedback, users actively tested the function with diverse inputs, including words and phrases in English and Chinese, abbreviations in pinyin, internet slang, and other language forms such as emoji, kaomoji, coded texts, etc. The findings highlight the importance of closer collaboration among computer scientists, translation scholars, and platform designers to better understand and improve translation technologies in real-world communicative contexts.
@@ 1 Introduction
Looking at Xiaohongshu’s development over the years, a recurring theme is its orientation towards international markets. The original app was known as ‘Hong Kong Shopping Guide’, targeting Chinese tourists seeking shopping recommendations outside mainland China [11]. In December 2013, six months after the company’s founding, Xiaohongshu launched a community dedicated to sharing overseas shopping experiences – an initiative that the company itself highlights as one of the key milestones in the early development [15]. Subsequent initiatives further reinforced this international vision. In May 2017, the company introduced ReDelivery, an international logistics service designed to facilitate cross-border e-commerce. Later that year in August, Xiaohongshu established its Global Technology Headquarters in Wuhan, China. In an interview, one of the founders emphasised their global vision and their ambition to position the platform as a space for sharing lifestyle experiences for Chinese and overseas users alike [1].
@@
As the platform gradually expanded its international reach, issues of cross-lingual communication became increasingly relevant. While Xiaohongshu had long been oriented primarily towards Chinese-speaking users, its global aspirations and cross-border commerce infrastructure created conditions in which multilingual interaction could become more prominent. This dynamic became particularly visible in 2025, when the platform experienced a sudden influx of international users.
@@
In January 2025, Xiaohongshu attracted a large number of international users following renewed discussions about a potential ban on TikTok in the United States. The literature shows that more than three million English-speaking users migrated to the platform [4], and the influx of these users propelled Xiaohongshu to become the most downloaded application on Apple’s US App Store on 13 January 2025 [14]. Many of these users identified themselves as ‘TikTok refugees’, and discussions surrounding this phenomenon on Xiaohongshu accumulated over 2.45 billion views within days [17].
@@
The emergence of this multilingual community revealed the limitations that language barriers imposed on cross-cultural communication on the platform. In response to these challenges, only a couple of days later since the influx, Xiaohongshu introduced a built-in translation feature on 18 January 2025, announced by an official post titled bilingually as ‘小红书翻译功能来啦!Translation is coming’. As any fluent bilingual speaker of Chinese and English would notice, let alone professional translators, this title is far from a good translation: the Chinese title, meaning ‘Xiaohongshu’s translation function is here’, is clear and informative, whereas the English text is ambiguous and confusing when read on its own.
@@
In this context, this study investigates user responses to the launch of Xiaohongshu’s built-in MT function. By analysing users’ reactions, comments, and the texts they used to try out the translation feature, the study examines how the new functionality was perceived and experimented by users within the community. Specifically, it addresses the following research questions:
@@ 2 Literature Review
MT has become an essential technological feature for facilitating multilingual interaction on social media platforms such as Instagram, TikTok, and X (formerly, Twitter). On these platforms, the MT systems are expected not only to translate textual content, but also support multilingual sense-making and interaction that allow users from different linguistic backgrounds to participate in shared communicative spaces [9]. However, idiosyncratic features of social media content present particular challenges for MT systems [3] and the increasing globalisation trend intensifies the demand for effective MT-assisted cross-lingual communication [6]. Despite the increasing availability of built-in translation functions on social media platforms, the effectiveness and social implications of these systems remain contested [8].
@@
One strand of literature emphasises the transformative potential of LLMs for global communication. For example, Yan et al. [16] argue that LLMs enable new forms of international interaction by dismantling barriers of language, culture, and geography. They cite Xiaohongshu as an example, stating that the platform’s LLM-driven translation tools allow users from diverse linguistic backgrounds to communicate ‘seamlessly’ in real time. Notably, such claims found in literature often rely on technological optimism and do not examine translation accuracy or user experiences in practice.
@@
Despite these emerging insights, a significant gap is evident in the literature. While technical studies focus on model development and performance, and social science research occasionally references translation as a supporting feature, there is little empirical investigation into how the translation function on Xiaohongshu was experienced by users when it was first launched, and the perceptions, interpretations, and practices of users engaging with the platform’s translation tools remain largely unexplored. Addressing this gap is especially important given the platform’s growing international user base and the increasing role of automated translation in shaping intercultural and interlingual communication within social media ecosystems.
@@ 3.1 Data Collection
Second, comments associated with these posts were retrieved using the open-source web crawling tool MediaCrawler. [^2] https://github.com/NanmiCoder/MediaCrawler For each of the 11 identified posts, the crawler was used to collect the top 300 first-level comments (as ranked by the platform’s default sorting and visible to generic users within the application interface), along with their corresponding sub-comments, which typically contained reflections on the translation of segments in the first-level comments or additional testing segments. This approach ensured that the dataset reflected the most visible and highly engaged user discussions surrounding the newly introduced feature.
@@ 3.2.2 Sentiment Analysis
To address these challenges and account for the complex linguistic and multimodal nature of the dataset, this study adopts a triangulated LLM-plus-human approach using two flagship models (GPT-5 mini [^3] https://developers.openai.com/api/docs/models/gpt-5-mini and DeepSeek-V3.2444https://huggingface.co/deepseek-ai/DeepSeek-V3.2) supplemented by human annotation of randomly selected samples following an aspect-based approach. Each comment was independently classified by the two models, and their outputs were compared to derive consensus results (n = 4519; 67.22%). To assess the reliability of the automated classifications, a randomly selected 10% subset (n = 452) of the consensus-labelled data, distributed across the four use cases, was manually annotated by a human coder. Agreement between the human annotations and the model consensus results was evaluated using Cohen’s kappa ($\kappa$ = .81), indicating a high level level of agreement. [^5] Human annotation revealed occasional misclassifications in the consensus-labelled data, particularly in comments containing only ‘hahahaha’ or slang expressions involving swearing words used to intensify excitement. As detailed analysis of these patterns lies beyond the scope of the present study, they are not examined further here. This mixed approach combines the scalability of automated analysis with human validation, helping to mitigate potential model bias and improve the robustness of the analytical results. Table 1 shows an overview of the collected and analysed data.
@@ 4.1 Sentiments
As shown in Figure 2, sentiment distributions varied across the four use cases. Comments associated with post content and comment translation constituted the largest share of the dataset (n = 2,403), with the highest proportion of positive sentiment (n = 1,935; 80.5%) across the four use cases and the lowest of negative (n = 370; 15.4%). For direct message translation (n = 144), while it hit the lowest for both positive (n = 84; 58.3%) and neutral (n = 5; 3.5%) comments, the proportion of negative comments (n = 55; 38.2%) was the highest. Similarly, despite bilingual posting (n = 51) showed a generally positive distribution (n = 34; 66.7%), it had the highest proportion of neutral comments (n = 9; 17.6%) and the negative responses (n = 8; 15.7%) were among the lower range. Bilingual subtitling (n = 447) showed a more balanced distribution, where the proportion of positive (n = 311) and negative (n = 115) comments reached 69.6% and 25.7% with neutral staying at 4.7% (n = 21).
@@
For positive comments, the most frequently occurring theme was on the translation function, where users commented favourably on the translation performance, especially for translating certain texts such as abbreviations and internet slangs (e.g., 网络热词也可以翻译 Hot internet expressions can be translated as well), or the usefulness of the feature in general (e.g., 好厉害的翻译 outstanding translation). Other positive themes included appreciation of the existence of the feature itself (e.g., 终于有翻译功能了 a translation function at last), references to the platform when praising the feature (e.g., 小红书超厉害 Xiaohongshu is super), and expressions of gratitude toward the developers or technical staff responsible for implementing the function (e.g., 给程序员加鸡腿 辛苦啦 give more chicken legs to the programmers thanks for the hardwork). A smaller number of comments also highlighted the speed with which the feature was introduced.
Neutral comments largely consisted of practical inquiries or suggestions related to the feature (e.g., 能不能做一个视频字幕带翻译功能呀 can you enable the translation function for subtitles in videos). Some of these inquiries were promptly addresses by the platform as can be seen in the positive comments. Some users asked questions about how the translation function works (e.g., 这是什么语言什么翻译原理 What language is this and what is the translation mechanism of this, while others proposed improvements (e.g., 建议再优化一下翻译功能 文字里面带表情图的没有翻译选项 The translation feature could be improved. Text embedded in images with emojis cannot currently be translated.). A small number of comments involved users sharing information about their device system or application version when explaining why they could not access the feature.
Negative themes primarily concerned technical issues or limitations of the feature. The most common negative theme was dysfunction, where users reported problems with the translation function not operating as expected (e.g., 我为啥看不到有翻译啊 why can’t I see a translation). Additional negative comments referred to translation quality or performance issues (e.g., 小红书怎么被翻译成了红笔记 how come Xiaohongshu was translated as Hongbiji), incompatibility with certain devices or application versions, and occasional remarks concerning staff welfare or working conditions.
@@ 4.2 Test Segments
The analysis on test segments inserted by users shows several recurring categories, as shown in Figure 4. The most common category consisted of natural language inputs, including texts in multiple languages such as English, Chinese, Cantonese, Chinese (Min), Burmese, Japanese, Russian, Malay, and Latin. In addition, users frequently experimented with language varieties and creoles, including Singlish (Singapore English) and Chinglish (Chinese English). In the English and Chinese segments, most entries consist of greeting (e.g., 外国朋友你们好 hello foreign friends), isolated lexical items (e.g., abandon, which often appears among the first words in English vocabulary textbooks), or proper nouns (e.g., KFC).
@@
A second group of test inputs involved artificial or styled language formats, including coding schemes such as American Standard Code for Information Interchange (ASCII) (e.g., ‘73 76 79 86 69 89 79 85’ was translated as ‘I LOVE YOU’), JavaScript, Braille, and Morse code, as well as kaomoji (e.g., /o_o\ was translated as 猫猫叹气, a meow sigh) and stylised codes (e.g., ‘●○●●○●●○●’ was translated as 星星之火可以燎原, a single spark can start a prairie fire – a wellknown saying in Chinese).
@@ 5 Discussion
5.1 Perceived Performance Across Use Cases
User responses suggest that perceptions of the translation feature were shaped strongly by the context in which it was used. Among the four use cases examined, translating posts and comments was generally perceived most positively. These functions were frequently described as useful for understanding foreign-language content and enabling cross-lingual interaction on the platform. By contrast, multimodal applications including bilingual posting and subtitling received comparatively less favourable reactions. Users reported more mixed experiences with these features, often pointing to inconsistencies or limitations in output quality and usability.
This variation indicates that user perceptions of translation quality are closely tied to the communicative context in which the function operates. Text-based translation of posts and comments appears to align more directly with users’ immediate communicative needs, while multimodal implementations introduce additional layers of complexity. The comparatively weaker reception of multimodal features may therefore reflect both technical challenges and differences in how users evaluate translation usefulness in everyday platform activities. The platform’s subsequent consolidation of some of these multimodal functions into a broader ‘content translation’ function embedded in the application settings suggests that early user feedback may have informed later design adjustments.
@@ 5.1 Perceived Performance Across Use Cases
Another dimension shaping perceived performance was the type of input users chose when testing the feature. A noticeable proportion of comments focused on the system’s ability to decode unconventional inputs such as kaomoji, stylised symbols, ASCII-coded text, Morse code, or other textual and/or visual encoded forms. Successful translations of these inputs often generated surprise and enthusiastic reactions, which users interpreted as evidence of the system’s technical sophistication. At the same time, relatively few tests involved complex sentences or extended exchanges in everyday languages. In many cases, tests of natural language translation only consisted of greetings, isolated words, or proper nouns. As a result, positive impressions of translation performance were frequently formed on the basis of these atypical or simplified inputs rather than on sustained multilingual communication.
5.2 Evaluation of Translation Function
@@
This pattern of experimentation highlights an important limitation in how users assess MT performance in this case study. When users test artificial codes to encode texts or symbolic forms to deliver intended meaning, they typically know the expected output and can therefore verify the correctness of the translation directly. In contrast, when translating everyday languages that they do not understand, users often lack the linguistic knowledge required to evaluate the accuracy of the output, not to mention the cultural knowledge to evaluate its appropriateness. Under such circumstances, the translation produced by the system may by default be accepted by users as proper equivalence.
@@
The popularity of testing unconventional segments therefore creates a potential risk in shaping user perceptions. Successful decoding of symbolic inputs may lead users to infer that the system performs equally well when translating natural languages used in everyday communication. However, translation in real-world contexts involves far more than decoding seemingly cool symbolic representations or finding the corresponding words and phrases in another language as the purpose that a dictionary would serve. It requires interpretation of meaning within specific social, cultural, and pragmatic contexts. Treating translation primarily as a technical decoding process may hence obscure the situated and interpretive nature of cross-lingual communication.
@@
These findings underscore the importance of raising public awareness about the limitations and complexities of automated translation. While the ability of LLMs to process diverse symbolic forms may appear impressive, such demonstrations do not necessarily reflect performance in everyday multilingual interactions. Encouraging more informed understandings of translation as a socially situated activity may therefore help users engage with MT outputs more critically.
5.3 Translation in a Prominent Socio-Technical System

### 03 Apr 2026 v2

@@ Abstract
This paper examines user reactions to the launch of the machine translation (MT) feature on Xiaohongshu, a Chinese social media and e-commerce platform, in January 2025. Drawing on a dataset of 6,723 comments collected from 11 official posts promoting the translation function, this paper combines sentiment analysis with thematic analysis to investigate how users perceived and experimented with the function. Results show that reactions were generally positive, although concerns regarding functionality, accessibility, and translation accuracy were also expressed. In addition, users actively tested the function with inputs that are atypical for everyday online communication, including stand-alone words and phrases, abbreviations, internet slang, and symbolic or encoded forms. Successful decoding of these texts elicited positive responses, while testing of more conventional language remained fairly limited. This could lead to uncritical acceptance of MT outputs by users, highlighting the importance of closer collaboration among computer scientists, translation scholars, and platform designers to improve MT performance and promote informed user engagement in real-world scenarios.
@@ 1 Introduction
Looking at Xiaohongshu’s development over the years, a recurring theme is its orientation towards international markets. The original app was known as ‘Hong Kong Shopping Guide’, targeting Chinese tourists seeking shopping recommendations outside mainland China [19]. In December 2013, six months after the company’s founding, Xiaohongshu launched a community dedicated to sharing overseas shopping experiences – an initiative that the company itself highlighted as one of the key milestones in its early development [24]. Subsequent initiatives further reinforced this international vision. In May 2017, the company introduced ReDelivery, an international logistics service designed to facilitate cross-border e-commerce. Later that year in August, Xiaohongshu established its Global Technology Headquarters in Wuhan, China. In an interview, one of the founders emphasised their global vision and their ambition to position the platform as a space for sharing lifestyle experiences for Chinese and overseas users alike [2].
@@
As the platform gradually expanded its international reach, issues of cross-lingual communication became increasingly relevant. While Xiaohongshu had long been oriented primarily towards Chinese-speaking users, its global aspirations and cross-border commerce infrastructure created conditions in which multilingual interaction could become more prominent. This dynamic saw a turning point in January 2025, when the platform experienced a sudden influx of international users.
@@
Following renewed discussions about a potential ban on TikTok in the United States in the beginning of 2025, Xiaohongshu attracted a large number of international users. The literature shows that more than three million English-speaking users migrated to the platform [8], and the influx of these users propelled Xiaohongshu to become the most downloaded application on Apple’s US App Store on 13 January 2025 [23]. Many of these users identified themselves as ‘TikTok refugees’, and discussions surrounding this phenomenon on Xiaohongshu accumulated over 2.45 billion views within days [26].
@@
The surge of this multilingual community revealed the limitations that language barriers imposed on cross-cultural communication on the platform. In response to these challenges, only a couple of days later since the influx, Xiaohongshu promptly introduced a machine translation (MT) function powered by large language models (LLMs) on 18 January 2025, announced by an official post titled bilingually as ‘小红书翻译功能来啦!Translation is coming’. As any fluent bilingual speaker of Chinese and English would notice, let alone professional translators, this title is far from a good translation: the Chinese title, meaning ‘Xiaohongshu’s translation function is here’, is clear and informative, whereas the English text can be ambiguous and confusing when read on its own. Following this post, a series of official posts were published over the subsequent week, introducing and promoting the MT feature across different use scenarios. These posts collectively formed an informational package, making users aware of the feature’s existence and use cases, such as translating posts, comments, and chats.
@@
Building on this launch context, this study investigates user responses to the launch of Xiaohongshu’s built-in MT function in its launch context. By analysing users’ reactions, comments, and the texts they used to try out the translation feature, the study examines how the new functionality was perceived and experimented by users within the community while also reflecting on the broader lessons that can be drawn from this pioneering attempt to move LLM-powered MT from laboratory settings into mass-user communication. Specifically, it addresses the following research questions:
@@ 2 Literature Review
MT has become an essential technological feature for facilitating multilingual interaction on social media platforms such as Instagram, TikTok, and X (formerly, Twitter). On these platforms, the MT systems are expected not only to translate textual content, but also support multilingual sense-making and interaction that allow users from different linguistic backgrounds to participate in shared communicative spaces [15]. However, idiosyncratic features of social media content present particular challenges for MT systems [6, 20] and the increasing globalisation trend intensifies the demand for effective MT-assisted cross-lingual communication [10]. Despite the increasing availability of built-in translation functions on social media platforms, the effectiveness and social implications of these systems remain contested [12].
@@
One strand of literature emphasises the transformative potential of LLMs for global communication. For example, Yan et al. [25] argue that LLMs enable new forms of international interaction by dismantling barriers of language and culture. They use Xiaohongshu as an example, stating that the platform’s LLM-driven translation tools allow users from diverse linguistic backgrounds to communicate ‘seamlessly’ in real time. Notably, such claims found in literature often rely on technological optimism and do not examine translation accuracy or user experiences in practice.
@@
Despite these emerging insights, a significant gap is evident in the literature: While technical studies focus on model development and performance, and social science research occasionally references translation as a supporting feature, there is little empirical investigation into how the translation function on Xiaohongshu was experienced by users when it was first launched, and the perceptions, interpretations, and practices of users engaging with the platform’s translation tools remain largely unexplored. Addressing this gap is especially important given the platform’s growing international user base and the increasing role of MT in shaping intercultural and interlingual communication within social media ecosystems.
@@ 3.1 Data Collection
Second, comments associated with these posts were retrieved using the open-source web crawling tool MediaCrawler. [^2] https://github.com/NanmiCoder/MediaCrawler For each of the 11 identified posts, the crawler was used to collect the top 300 first-level comments (as ranked by the platform’s default sorting and visible to generic users within the application interface), along with their sub-comments, which typically contained reflections on the translation of segments in the first-level comments or additional testing segments. This approach ensured that the dataset reflected the most visible and highly engaged user discussions on the newly introduced feature.
@@ 3.2.2 Sentiment Analysis
To address these challenges and account for the complex linguistic and multimodal nature of the dataset, this study adopts a triangulated LLM-plus-human approach using two flagship models (GPT-5 mini [^3] https://developers.openai.com/api/docs/models/gpt-5-mini and DeepSeek-V3.2444https://huggingface.co/deepseek-ai/DeepSeek-V3.2) supplemented by human annotation of randomly selected samples following an aspect-based approach. Each comment was independently classified by the two models, and their outputs were compared to derive consensus results (n = 4519; 67.22%). To assess the reliability of the automated classifications, a randomly selected 10% subset (n = 452) of the consensus-labelled data, distributed across the four use cases, was manually annotated by a human coder. Agreement between the human annotations and the model consensus results was evaluated using Cohen’s kappa (κ = .81), indicating a high level level of agreement. [^5] Human annotation revealed occasional misclassifications in the consensus-labelled data, particularly in comments containing only ‘hahahaha’ or slang expressions involving swearing words used to intensify excitement. As detailed analysis of these patterns lies beyond the scope of the present study, they are not examined further here. This mixed approach combines the scalability of automated analysis with human validation, helping to mitigate potential model bias and improve the robustness of the analytical results. Table 1 shows an overview of the collected and analysed data across the four use cases.
@@ 4.1 Sentiments
As shown in Figure 2, sentiment distributions varied across the four use cases. Comments associated with post content and comment translation constituted the largest share of the dataset (n = 2,403), with the highest proportion of positive sentiment (n = 1,935; 80.5%) across the four use cases and the lowest of negative (n = 370; 15.4%). For direct message translation (n = 144), as it hit the lowest for both positive (n = 84; 58.3%) and neutral (n = 5; 3.5%), the proportion of negative comments (n = 55; 38.2%) was unsurprisingly the highest.For bilingual posting (n = 51), although it showed a generally positive distribution (n = 34; 66.7%), it had the highest proportion of neutral comments (n = 9; 17.6%) and the negative responses (n = 8; 15.7%) were at the lower range. Bilingual subtitling (n = 447) showed a more balanced distribution, where the proportion of positive (n = 311) and negative (n = 115) comments reached 69.6% and 25.7% with neutral staying at 4.7% (n = 21).
@@
For positive comments, the most frequently occurring theme was on the translation function, where users commented favourably on the translation performance, especially for translating certain texts such as abbreviations and internet slangs (e.g., 网络热词也可以翻译 it can translate popular internet expressions as well), or the usefulness of the feature in general (e.g., 好厉害的翻译 outstanding translation). Other positive themes included appreciation of the existence of the feature itself (e.g., 终于有翻译功能了 a translation function at last), references to the platform when praising the feature (e.g., 小红书超厉害 Xiaohongshu is super), and expressions of gratitude towards the developers or technical staff responsible for implementing the function (e.g., 给程序员加鸡腿 辛苦啦 give more chicken legs to the programmers; thanks for the hard-work). A smaller number of comments also highlighted the speed with which the feature was introduced.
Neutral comments largely consisted of practical inquiries or suggestions related to the feature (e.g., 能不能做一个视频字幕带翻译功能呀 can you enable the translation function for subtitles in videos). Some inquiries of this kind were promptly addresses by the platform, as can be seen in the positive comments above. In addition, some users asked questions about how the translation function works (e.g., 这是什么语言什么翻译原理 What language is this and what is the translation mechanism of this), while others proposed improvements (e.g., 建议再优化一下翻译功能 文字里面带表情图的没有翻译选项 The translation feature could be improved. Text embedded in images with emojis don’t have the translation option). A small number of comments involved users sharing information about their device system or application version when explaining why they could not access the feature.
Negative themes primarily concerned technical issues or limitations of the feature. The most common negative theme was dysfunction, where users reported problems with the translation function not operating as expected (e.g., 我为啥看不到有翻译啊 why can’t I see the translation). Additional negative comments referred to translation quality or performance issues (e.g., 小红书怎么被翻译成了红笔记 how come Xiaohongshu was translated as Hongbiji), incompatibility with certain devices or application versions, and occasional remarks concerning staff welfare or working conditions.
@@ 4.2 Test Segments
The analysis on test segments posted by users shows several recurring categories, as shown in Figure 4. The most common category consisted of natural language inputs, including texts in multiple languages such as English, Chinese, Cantonese, Chinese (Min), Burmese, Japanese, Russian, Malay, and Latin. In addition, users frequently experimented with language varieties and creoles, including Singlish (Singapore English) and Chinglish (Chinese English). In the English and Chinese segments, most entries consist of greeting (e.g., 外国朋友你们好 hello foreign friends), isolated lexical items (e.g., abandon – which often appears among the first words in English vocabulary textbooks), or proper nouns (e.g., KFC).
@@
A second group of test inputs involved artificial or styled language formats, echoing the content of one of the official posts P04 that featured a string of coded text in its title: 请翻译:[symbols] (Please translate: [symbols]). This group included coding schemes such as American Standard Code for Information Interchange (ASCII) (e.g., ‘73 76 79 86 69 89 79 85’ was translated as ‘I LOVE YOU’), JavaScript, Braille, and Morse code, as well as kaomoji (e.g., [kaomoji] was translated as 猫猫叹气, a meow sigh) and stylised codes (e.g., [symbol code] was translated as 星星之火可以燎原, a single spark can start a prairie fire – a well-known saying in Chinese).
@@ 5 Discussion
5.1 User Feedback to the New Function
The built-in MT function on Xiaohongshu represents a pioneering effort to deploy LLM-based MT on a mass-user platform outside the lab environment, providing valuable insights into both successes and shortcomings. In terms of its basic configuration, the built-in function was activated based on users’ system language settings. For users whose system language was set to a language other than Chinese, the translation function was automatically enabled for Chinese content on the platform. Conversely, for users with Chinese as their system language, the function became available when non-Chinese content was shown. Yet, this information was not communicated clearly in the official posts – only ‘click the translate tab’ and ‘update to the newest version’ were mentioned, and when it was tested in multilingual settings, problems started to emerged, with users complaining not seeing translations they expected or not seeing the translation option at all.
User responses further suggest that perceptions of the translation feature differed across use cases, indicating perceptions of MT functionality were shaped by the context in which it was used. Among the four use cases examined, translating posts and comments was generally perceived most positively. By contrast, multimodal applications including bilingual posting and subtitling received comparatively less favourable reactions. The comparatively weaker reception of multimodal features reflects how contextualised interface design shaped users’ evaluation of the MT function. Text-based translation of posts and comments was supported by clear signposting of a clickable ‘translate’ option next to translatable content, which fits users’ intuitive expectations of how the function should work. In contrast, accessing translation in chats requires users to long-press the text before the ‘translate’ option appears. Both interaction mechanism were clearly presented in the official posts but the key difference was with the interface design. These multimodal use cases seem to have introduced additional layers of complexity, compounded by the absence of efficient onboarding information for users, as reflected in the negative comments on usability and accessibility.
The platform’s subsequent consolidation of language settings and multimodal functions into a broader ‘content translation’ function embedded in the application settings suggests that early user feedback may have informed later design adjustments in a retrospective manner. Reflecting on this trajectory of updates, it becomes evident that launching an MT tool requires careful planning that extends beyond technical model performance. Early deployment should incorporate a comprehensive launch strategy that accounts for user experience, ideally validated through pilot studies, rather than moving directly from a lab setting with limited stakeholders to full-scale public release. While practical pressures such as time constraints and business priorities inevitably influence deployment decisions, considerations of quality and usability should remain central when releasing an MT tool to a mass audience. Such an approach can be guided by emerging research-based, evidence-driven ethical frameworks [5], ensuring reliability, accessibility, and trustworthiness for all users.
5.2 Evaluation of Translation Quality
@@
Another key theme emerged from the type of source texts that users chose when testing the feature. A noticeable proportion of comments focused on the system’s ability to decode unconventional inputs in textually and/or visually encoded forms.  [^7] This phenomenon could reasonably be linked to the official post P04 that presented a string of symbols as the source text for MT and invited encoded texts from users to test the translation function (if true, this might indicate yet another flaw in the launch package discussed in Section 5.1). The results show that the successful translations of these inputs often generated enthusiastic reactions, which users interpreted as evidence of the MT system’s advanced performance. At the same time, relatively few tests involved complex sentences or extended exchanges in everyday languages. As a result, positive impressions of translation performance were frequently formed on the basis of these atypical or simplified inputs rather than on sustained multilingual communication.
@@
Source texts that users experimented with for the MT feature were very different from the source texts proposed by Xiaohongshu users for language learning purposes as report in Chen [7]. When testing artificial codes to encode texts or symbolic forms to deliver intended meaning, users typically know the expected output and can therefore verify the correctness of the translation directly. In contrast, when translating everyday languages that they understand very little, users often lack the linguistic knowledge or even confidence required to evaluate the accuracy of the output, not to mention the cultural knowledge to evaluate its appropriateness. This is in line with Qiu and Pym [18], where MT errors evidenced language learners’ lack of self-confidence by trusting the (mis-)translated subtitles more than themselves. For users’ engagement with MT outputs examined in the current study, it is more likely to be the case where users were not confident in their second language at all, if there is any. The absence of linguistic knowledge as such can potentially lead to blind trust in MT outputs and the translation produced by the built-in system may by default be accepted by users as accurate and appropriate.
@@
The popularity of testing unconventional segments therefore creates a potential risk in shaping user perceptions and experience. Successful decoding of symbolic inputs may lead users to infer that the system performs equally well when translating natural languages used in everyday communication. However, translation in real-world contexts involves far more than decoding seemingly cool symbolic representations or finding the corresponding words and phrases in another language as the purpose that a dictionary would serve. It requires interpretation of meaning within specific social, cultural, and pragmatic contexts. Treating translation primarily as a technical decoding process may hence obscure the situated and interpretive nature of cross-lingual communication, not to mention the baseline quality of MT outputs deployed on other Chinese social network services such as Wechat is not without problems [16].
@@
These findings underscore the importance of the ongoing effort for improving MT literacy – not only among translation students and educators, but also professional translators and other stakeholders – and for raising public awareness about the limitations and complexities of MT [1, 14, 3]. Encouraging more informed understandings of translation as a socially situated activity may therefore help users engage with MT outputs more critically.
5.3 Translation in a Socio-Technical System